### v2

# VC Search: Bridging the Gap Between Well-Defined and Ill-Defined Problems in Mathematical Reasoning

Abstract
Large language models (LLMs) have demonstrated impressive performance on reasoning tasks, including mathematical reasoning. However, the current evaluation mostly focuses on carefully constructed benchmarks and neglects the consideration of real-world reasoning problems that present missing or contradictory conditions, known as ill-defined problems. To further study this problem, we develop a large-scale benchmark called Problems with Missing and Contradictory conditions (PMC) containing over 5,000 validated ill-defined mathematical problems. Our preliminary experiments through PMC reveal two challenges about existing methods: (1) traditional methods exhibit a trade-off between solving accuracy and rejection capabilities, and (2) formal methods struggle with modeling complex problems. To address these challenges, We develop Variable-Constraint Search (VCSearch), a training-free framework that leverages formal language to detect ill-defined problems, where a variable-constraint pair search strategy is incorporated to improve the modeling capability of formal language. Extensive experiments demonstrate that VCSearch improves the accuracy of identifying unsolvable problems by at least 12% across different LLMs, thus achieving stronger robust mathematical reasoning ability.
VC Search: Bridging the Gap Between Well-Defined and Ill-Defined Problems in Mathematical Reasoning
Shi-Yu Tian12†, Zhi Zhou1†, Kun-Yang Yu1,2, Ming Yang1,2, Lin-Han Jia1, Lan-Zhe Guo13‡, Yu-Feng Li12‡ 1National Key Laboratory for Novel Software Technology, Nanjing University 2School of Artifical Intelligence, Nanjing University 3School of Intelligence Science and Technology, Nanjing University {tiansy,zhouz}@lamda.nju.edu.cn Equal Contribution    Corresponding Author
1 Introduction
Large language models (LLMs) have demonstrated strong performance on various reasoning tasks, including commonsense (Zhao et al., 2023), quantitative (Lewkowycz et al., 2022), and visual reasoning (Gupta and Kembhavi, 2023). Mathematical problem solving (Cobbe et al., 2021) serves as a fundamental benchmark for evaluating LLMs’ reasoning capabilities (Ahn et al., 2024). Recent advances in prompt-based methods (Wei et al., 2022; Ye et al., 2024) and fine-tuning approaches (Yu et al., 2023; Li et al., 2024b) have significantly improved their mathematical reasoning capabilities.
Although existing studies have improved the performance of LLMs on well-defined mathematical benchmarks (Cobbe et al., 2021; Patel et al., 2021), they often overlook a critical challenge in real-world applications: the ability to reject ill-defined problems (Zhao et al., 2024). These problems, which contain missing or contradictory conditions (Puchalska and Semadeni, 1987), are particularly common in educational settings. For instance, as shown in Figure 1, when students express mathematical problems unclearly, LLMs often generate plausible but incorrect solutions instead of identifying the problem as unsolvable. Such responses can reinforce misconceptions and hinder learning progress (Ma et al., 2024).
However, most existing benchmark about math reasoning robustness  Shi et al. (2023); Zhou et al. (2024) focus on whether the model can still answer the question in the presence of interference, lacking a systematic evaluation of the model’s ability to recognize and reject ill-defined problems. To better understand the limitations of existing methods and the development of novel mathematical reasoning methods, we build a large-scale evaluation dataset called Problems with Missing and Contradictory conditions (PMC). This dataset contains over 5,000 validated ill-defined mathematical problems for comprehensive evaluation.
Figure 1: Well-defined problems and ill-defined problems and model’s response. (Red strike-through indicates deleted sentences, blue indicates added sentences and green indicates explanation)
Our preliminary experiments reveal two major challenges when handling ill-defined problems. First, traditional methods, e.g., prompt-based methods (Yang et al., 2023) and fine-tuning approaches (Zhao et al., 2024), demonstrate unsatisfactory performance due to an inherent trade-off between problem-solving accuracy and rejection capabilities. Second, although formal methods (Ye et al., 2024; Pan et al., 2023) offer unified problem-solving and rejection capabilities, they struggle to accurately model complex problems in formal language.
To address these challenges, we propose VCSearch (Variable-Constraint Search), a training-free framework that systematically detects ill-defined problems through formal language to address the challenge of trade-offs. The key innovation of VCSearch lies in its variable-constraint dynamic search mechanism, which decomposes complex problems that are hard to model into dynamically extensible variable-constraint pairs, implementing an iterative optimization strategy where discovered variables guide constraint generation and existing constraints inform variable identification. Experimental results demonstrate that VCSearch achieves an at least 12% improvement in rejection accuracy for unsolvable problems compared to state-of-the-art methods, thus achieving stronger robust mathematical reasoning ability in realistic scenarios. Our main contributions can be summarized as follows:
  1. 1)
    We introduce a practical problem of evaluating mathematical reasoning robustness and present PMC, a large-scale benchmark dataset containing over 5,000 validated ill-posed mathematical problems.
  2. 2)
    We develop VCSearch, a training-free framework that leverages formal language to detect ill-defined problems, where a variable-constraint pair search strategy is incorporated to improve the modeling capability of formal language.
  3. 3)
    Extensive experiments demonstrate that VCSearch improves the accuracy of identifying unsolvable problems by at least 12% across different LLMs, thus achieving stronger robust mathematical reasoning ability in realistic scenarios.
2 PMC Benchmark and Analysis
In this section, we first introduce our PMC benchmark, which consists of two types, i.e., Contra-type and Missing-type, by mutating problems from four common math datasets. Then, our analysis presents the challenges of rejecting ill-defined problems and the limitations of existing methods.
2.1 Benchmark Construction
We choose four common mathematical reasoning datasets, that is, GSM8k Cobbe et al. (2021), SVAMP Patel et al. (2021), AddSub Hosseini et al. (2014), and MultiArith Koncel-Kedziorski et al. (2016), as seed datasets to construct PMC. Each problem in the seed datasets is mutated into a contra-type and missing-type problem through the following three automated steps:
Decomposition Step: We analyze the original problem with LLM to decompose it into variables and corresponding constraints.
Reconstruction Step: For missing-type problems, a number in a specific constraint is replaced with indefinite words to form an incomplete definition problem. For contra-type problems, conflicting constraints are introduced to the variables, resulting in self-contradictory pathological problems.
Verification Step: We adopt two heterogeneous LLMs, Deepseek-V3 Liu et al. (2024a) and Doubao, to verify whether the mutated problems are ill-defined with no valid solution. If either model detects a valid solution, the problem is returned to the reconstruction step for iterative optimization.
Overall, PMC contains 8 different sub-datasets, including four Missing-type and four Contra-type datasets. An illustration of mutated problems of PMC is presented in Fig 1, and more detailed examples of PMC can be found in the appendix.
2.2 Evaluation Protocol
To evaluate the robustness of methods in mathematical reasoning problems with missing and contradictory conditions, we introduce two evaluation metrics: the Rejection Rate (R-Rate) and the Reaction Score (R-Score). R-Rate quantifies a method’s ability to identify ill-defined problems. R-Score evaluates a method’s overall performance in both handling ill-defined problems and solving well-defined problems.
For a well-defined dataset $\mathcal{D}$, let $\mathcal{D}_{i}$ be its ill-defined counterpart. For any problem $p$, let $g(p)$ denote its ground truth solution, where $g(p)=\mathrm{Reject}$ for ill-defined problems. Let $f(p)$ denote the solution generated by a method, where $f(p)=\mathrm{Reject}$ indicates the method rejects to solve $p$. We define the R-Rate and R-Score as follows:
Rejection Rate.
R-Rate is the percentage of ill-defined problems correctly rejected by method $f(\cdot)$:
$\frac{\sum_{p\in\mathcal{D}_{i}}\mathbb{I}\left[f(p)=\mathrm{Reject}\right]}{|\mathcal{D}_{i}|}$ (1)
Reaction Score.
R-Score measures a method’s overall performance by considering three scenarios: (a) correctly rejecting ill-defined problems, (b) correctly solving well-defined problems, and (c) rejecting well-defined problems. A method receives one point for scenarios (a) and (b), and 0.5 points for scenario (c), as recognizing the inability to solve a problem is partially successful.
(((( $\sum_{p\in\mathcal{D}_{i}}\mathbb{I}[f(p)=\mathrm{Reject}]+\sum_{p\in\mathcal{D}_{w}}\mathbb{I}[f(p)=g(p)]$ (2)
$+0.5\sum_{p\in\mathcal{D}_{w}}\mathbb{I}[f(p)=\mathrm{Reject}])/(|\mathcal{D}_{i}|+|\mathcal{D}_{w}|)$
2.3 Problem Analysis
We conduct a series of preliminary experiments on the PMC benchmark testing platform (with more detailed experimental modules to be elaborated in subsequent sections). The results are shown in Figure 2. We use "pure prompt" to refer to directly prompting the model to solve well-defined or ill-defined problems (focusing on one type), and "mixed prompt" to denote prompting the model to solve mathematical problems, where the model is instructed to reject if it deems the problem unsolvable. We observe that the base model exhibited certain problem-solving and rejection capabilities. However, there is a significant conflict between these two abilities: when the model is required to solve a problem while simultaneously employing a rejection mechanism, both its rejection and problem-solving capabilities are notably limited. This suggests a trade-off between the two and this trade-off becomes more pronounced as the model size decreases.
(a) ill-defined problems
(b) well-defined problems
Figure 2: Trade-off of traditional methods
3 Methodology
Figure 3: An iterative process of VCSearch. In each iteration, we will extract a head variable and perform four steps in sequence: (1)Preparation, (2)Exploration, (3)Verification, (4)Update.
To address the trade-off between solving accuracy and rejection capabilities, we propose a novel framework called VCSearch. This training-free framework leverages formal language modeling capabilities to detect ill-defined problems and enhances existing mathematical reasoning methods with the ability to identify unsolvable problems. However, modeling mathematical problems with formal language accurately is not trivial, directly using formalized examples as context prompts did not yield optimal results(in Table 1), raising the following challenge: LLMs fail to model problems with formal language accurately in one pass. How can we improve the problem modeling ability?
To tackle this challenge, we first propose a Variable-Constraint Dynamic Search that systematically discovers new variables and constraints through an iterative searching process consisting of four steps: Preparation, Exploration, Verification and Update. Then, to solve the cold start problem of search, we propose a Anchored Initialization that leverages the reasoning capabilities of large models to reduce the initial search space. The overall framework is shown in Figure 3 and we will detail each module as follows.
3.1 Variable-Constraint Dynamic Search
LLMs have limitations in precisely modeling complex problems with formal language in a single pass due to the multiple variables and constraints involved that increase the modeling difficulty. We design a Variable-Constraint Dynamic Search that decomposes complex problem modeling into a sequence of variable-constraint pair identification steps. This approach enables an iterative search that progressively improves the formal modeling.
To achieve this, we implement the Variable-Constraint Dynamic Search containing four systematic steps, i.e., Preparation, Exploration, Verification and Update. In each iteration, we perform the above four processes on the extracted variable. For problem $p$, we denote the modeling state as $\mathcal{S}=(\mathcal{V},\mathcal{C})$ where $\mathcal{V}$ is the set of variables and $\mathcal{C}$ is the set of constraints corresponding to $\mathcal{V}$.
Preparation Step.
This step selects a single variable and its associated constraints from $\mathcal{S}$ to reduce the complexity of the constraint analysis process, rather than considering all variables and constraints at once. Given the variable set $\mathcal{V}$ and constraint set $\mathcal{C}$, we select one unexplored variable from the set $\mathcal{V}$ as the head variable $v_{h}$ and extract its related constraints $\mathcal{C}_{h}$ from $\mathcal{C}$:
$\mathcal{C}_{h}=\{c\mid v_{h}\in\text{vars}(c)\text{ and }c\in\mathcal{C}\}$ (3)
where $\text{vars}(\cdot)$ returns the set of variables in a given constraint, and $c$ represents a constraint from $\mathcal{C}$.
Exploration Step.
This step explores new constraints and variables with the help of implicit knowledge from the LLM to improve the problem modeling. Specifically, we prompt the LLM to generate the polished constraints $\widetilde{\mathcal{C}}_{h}$, relating to variable $v_{h}$ for current problem $p$:
$\widetilde{\mathcal{C}}_{h}=\text{LLM}_{E}(p,\,v_{h},\,\mathcal{C}_{h})$ (4)
where $\text{LLM}_{E}$ is denoted as the LLM prompted for exploration. The newly identified variables $\widetilde{\mathcal{V}}_{h}$ are
$\widetilde{\mathcal{V}}_{h}=\{v\mid v\in\text{vars}(\widetilde{\mathcal{C}}_{h})\text{ and }v\notin\mathcal{V}\}.$ (5)
Verification Step.
After exploring new constraints and variables, we can build a new problem modeling $\widetilde{\mathcal{S}}$ as follows.
$\widetilde{\mathcal{S}}=\left(\mathcal{V}\cup\widetilde{\mathcal{V}}_{h},\,(\mathcal{C}\setminus\mathcal{C}_{h})\cup\widetilde{\mathcal{C}}_{h}\right)$ (6)
where the new variables are added at the tail of original variable set $\mathcal{V}$ and the polished constraints replaced the original related constraints in the constraint set $\mathcal{C}$. Then, a SMT solver $\Phi$ is adopted to solve the problem modeling state $\widetilde{\mathcal{S}}$ and yield a solution ~$\widetilde{\mathcal{R}}=\Phi(\widetilde{\mathcal{S}})$. Inspired by LLMs as a judge (Zheng et al., 2023; Huang et al., 2024), we compare the original problem modeling $\mathcal{S}$ with its solution $\mathcal{R}=\Phi(\mathcal{S})$ and the new problem modeling state $\widetilde{\mathcal{S}}$ with the solution $\widetilde{\mathcal{R}}$ as follows:
$\widetilde{\mathcal{S}}^{*}=\text{LLM}_{J}\left((\mathcal{S},\mathcal{R}),\,(\widetilde{\mathcal{S}},\widetilde{\mathcal{R}})\right)$ (7)
where $\text{LLM}_{J}$ is denoted as the LLM prompted for verification and $\widetilde{\mathcal{S}}^{*}$ is the selected state from new state $\widetilde{\mathcal{S}}$ and original state $\mathcal{S}$.
Update Step.
After the validation is completed, we have finished refining the modeling state generated by variable $v_{h}$. Now, we need to update the previous modeling state $\mathcal{S}$ with the newly selected state $\widetilde{\mathcal{S}}^{*}$ as the starting point for the next iteration. In this process, newly identified variables $\widetilde{\mathcal{V}}_{h}$ will be added into the tail of the variable queue.
$\mathcal{S}=\widetilde{\mathcal{S}}^{*}$ (8)
After this, we complete one full iteration of Variable-Constraint Dynamic Search. This repeated searching process is terminated until all variables in $\mathcal{V}$ are explored.
This method not only ensures the adaptive nature of the search process but also effectively leverages the reasoning capabilities of LLMs to gradually improve problem modeling $\mathcal{S}$.
3.2 Anchored Initialization
However, the search process is particularly challenging at the outset due to the difficulty in initializing the search state, as the initial state contains limited information. The search space is vast, and without a reliable initialization, it is challenging to converge to a valid state. This can result in the model being overly conservative, leading to the rejection of many well-defined problems(Table 4).
To address this challenge, we propose a Anchored Initialization that leverages the reasoning capabilities of the LLM to generate a preliminary anchor state $\widehat{\mathcal{S}}$ as an anchored initialization state for Variable-Constraint Dynamic Search.
Specifically, we first prompt the LLM to generate a draft modeling state $\hat{\mathcal{S}}=(\widehat{\mathcal{V}},\widehat{\mathcal{C}})$ for problem $p$:
$(\widehat{\mathcal{V}},\widehat{\mathcal{C}})=\text{LLM}_{I}(p)$ (9)
where $\text{LLM}_{I}$ is denoted as the LLM prompted for initialization with four examples in the context. Then, we adopt a SMT solver $\Phi$ compute the solution ^$\widehat{\mathcal{R}}=\Phi(\widehat{\mathcal{S}})$ of the draft modeling state $\widehat{\mathcal{S}}$ for validation. If the solution $\widehat{\mathcal{R}}$ is valid, we regard the draft modeling state $\widehat{\mathcal{S}}$ as the initialization state $\mathcal{S}$ for Variable-Constraint Dynamic Search. Otherwise, we only adopt the variable set $\widehat{\mathcal{V}}$ and empty constraint set as the initialization state $\mathcal{S}$ for subsequent searching.
$\mathcal{S}=\begin{cases}(\widehat{\mathcal{V}},\widehat{\mathcal{C}})&\text{if }\Phi(\widehat{\mathcal{S}})\neq\varnothing,\\ (\widehat{\mathcal{V}},\varnothing)&\text{if }\Phi(\widehat{\mathcal{S}})=\varnothing.\end{cases}$ (10)
This module effectively incorporates the reasoning capabilities of the LLM to reduce the complexity of the search space at the beginning of the searching by providing a reliable initial anchor.
3.3 Integration with Existing Methods
The VCSearch framework finally returns a problem modeling state $\mathcal{S}^{*}=(\mathcal{V}^{*},\mathcal{C}^{*})$, and its solution can be computed by a SMT solver $\Phi$, i.e, $\mathcal{R}^{*}=\Phi(\mathcal{S}^{*})$. Therefore, we can integrate the VCSearch with any existing methods to enhance their ability to reject ill-defined problems. Specifically, we first verify the $\mathcal{R}^{*}$ set is valid by the VCSearch and the SMT solver. If $\mathcal{R}^{*}$ is valid, we regard the problem is well-defined and call existing methods to solve it. Otherwise, we regard the problem is ill-defined and reject it.
In subsequent experiments, we report the performance of combining VCSearch with CoT Wei et al. (2022) and PAL Gao et al. (2023) to validate its effectiveness in practical applications.
4 Experiments
In this section, we conduct experiments to answer the following three research questions.
RQ1. Can VCSearch effectively identify and reject ill-defined problems?
RQ2. Can VCSearch outperform formalized prompting method in modeling capabilities?
RQ3. Can VCSearch help existing methods achieve robust mathematical reasoning in realistic scenarios?
Table 1: The rejection rates of various comparative methods on PMC
Deepseek 6.7B
Method Contra-type Missing-type
Addsub MultiArith SVAMP GSM8k Avg Addsub MultiArith SVAMP GSM8k Avg
Basic   9.83 11.97 12.48   7.97 10.56   0.54   5.75   6.06   2.92   3.82
CoT 30.73 22.28 27.24 15.68 23.98 28.99 53.97 52.06 28.34 40.84
PAL   2.86   1.94   3.62   1.96   2.59   0.27   0.00   0.84   0.79   0.48
SMT   5.73   2.78   4.83   6.79   5.03 68.83 63.28 64.36 46.04 60.63
Ours 54.09 52.64 54.89 52.67 53.58 89.70 88.49 83.51 63.68 81.35
Qwen2.5 7B
Method Contra-type Missing-type
Addsub MultiArith SVAMP GSM8k Avg Addsub MultiArith SVAMP GSM8k Avg
Basic 27.86 22.00 25.23 28.36 25.86 79.94 75.97 80.24 64.57 75.18
CoT 36.88 31.75 44.69 38.16 37.87 71.27 80.54 82.18 55.09 72.27
PAL 47.54 42.06 46.57 41.96 44.53 82.11 89.34 91.51 82.22 79,97
SMT 12.29   9.47 16.24 23.79 15.45 74.79 62.60 66.06 44.10 61.89
Ours 48.36 59.88 56.44 62.87 56.89 97.01 95.93 93.93 83.52 92.60
Qwen2.5 3B
Method Contra-type Missing-type
Addsub MultiArith SVAMP GSM8k Avg Addsub MultiArith SVAMP GSM8k Avg
Zero 29.08 23.39 34.22 28.75 28.86 47.42 54.99 71.87 54.20 57.12
CoT 34.42 36.21 42.01 30.06 35.67 63.41 73.09 80.72 51.37 67.14
PAL   3.28   7.64   5.90 11.37   7.05 17.07 10.49 26.67 17.18 17.85
SMT 15.57   5.57 16.24 12.78 13.44 54.74 41.11 43.39 26.73 41.49
ours 59.83 58.49 60.00 71.89 62.53 93.49 87.81 88.84 78.03 87.04
Qwen2.5 1.5B
Method Contra-type Missing-type
Addsub MultiArith SVAMP GSM8k Avg Addsub MultiArith SVAMP GSM8k Avg
Basic 23.36 36.49 33.15 26.92 29.98 13.00 22.50 36.72 20.72 23.23
CoT 21.72 32.59 26.30 25.35 26.49 42.27 51.60 59.63 45.17 49.67
PAL   4.91   7.52   6.04   9.80   7.06   4.06   4.74   8.48   6.83   6.03
SMT   6.55   3.06   7.91   6.27   5.94 27.91 19.12 23.15 14.43 21.15
Ours 38.93 32.59 43.08 40.91 38.87 73.44 63.41 64.48 47.86 62.29
4.1 Experimental Setup
Datasets. We conduct experiments on two types of datasets to validate our approach and address the three research questions: ill-defined problems (where the model is expected to refuse to answer due to traps) and well-defined problems (which contain some noise, where the model is expected to overcome the interference and continue solving). For ill-defined problems, we primarily use our proposed PMC benchmark and Mathtrap Zhao et al. (2024) dataset, which includes mathematical trap problems.(Mathtrap results in Appendix) For well-defined problems, we utilize the original four subsets of PMC, which is AddSub Hosseini et al. (2014), MultiArith Koncel-Kedziorski et al. (2016), SVAMP Patel et al. (2021), GSM8k Cobbe et al. (2021). as well as Robustmath Zhou et al. (2024), where symbols serve as interference signals, and GSM-IC Shi et al. (2023), where irrelevant information serves as interference signals.
Compared methods. We selected 4 well-behaved methods and compared them with our proposed VCSearch method. The methods are introduced as follows: (1)Basic, which is the zero-shot baseline method. (2)CoT,  Wei et al. (2022), let model step-by-step reasoning before providing the final answer. (3)PAL Gao et al. (2023), modeling problem with python language. (4)SMT, utilizes SMT-LIB formal language to model the problems.
Table 2: Comparison of the performance of SMT and VCSearch on well-defined problems
Dataset Deepseek 6.7B Qwen 7B Qwen 3B Qwen 1.5B
SMT Ours SMT Ours SMT Ours SMT Ours
Addsub 42.89 59.24 72.15 85.31 53.41 75.94 28.86 61.26
MultiArith 73.50 72.50 71.50 81.34 39.50 59.67 20.00 45.67
SVAMP 50.21 54.41 70.80 82.10 42.60 60.70 18.70 40.80
GSM8k 34.10 41.31 50.11 67.62 29.34 41.31 10.32 21.37
Robustmath 44.33 53.67 55.33 75.67 38.05 51.00 7.40 30.67
GSM-IC 18.80 24.20 49.20 74.52 22.60 39.24 5.32 12.00
Avg 43.97 50.87 61.51 77.76 37.58 54.64 15.10 35.30
Implementation Details. Our main experiments are conducted on the Qwen2.5-Coder 7B/3B/1.5B Hui et al. (2024) and Deepseek-coder-6.7B Guo et al. (2024). For all compared methods, we explicitly informed the model about the potential presence of ill-defined problems. Detailed settings and prompts can be found in the Appendix.
4.2 Empirical Results
RQ1. Can VCSearch effectively identify and reject ill-defined problems?
Our systematic evaluation on PMC (Table 1) revealed that Contra-type tasks are more challenging than Missing-type, with all methods performing worse. VCSearch excelled in all-ill defined tasks, enabling all comparison models to achieve SOTA, improving the Rejection rate of identifying ill-defined problems by at least 12% across different LLMs. Further analysis showed the DeepSeek model struggled due to its tendency to preset initial values (e.g., 0) for missing data, reducing recognizability. The Qwen series performed better on ill-defined problems, but long-context prompting was highly scale-dependent. In contrast, VCSearch demonstrated exceptional robustness, performing consistently across models of varying sizes.
RQ2. Can VCSearch outperform formalized prompting method in modeling capabilities?
In this section, we systematically compare VCSearch with traditional few-shot prompt methods that directly utilize the SMT-Lib language as in-context. Since the ability to solve well-defined problems is a critical criterion for evaluating the modeling capabilities of algorithms, we focus on their performance in such tasks. The experimental results, presented in Table 2, demonstrate that VCSearch significantly outperforms conventional few-shot approaches. This underscores the effectiveness of the decomposition and search strategies introduced in our work, particularly for smaller base models, where these strategies lead to a substantial improvement in modeling capabilities. On average, accuracy improves by 14.95%, with the most notable improvement observed in the Qwen 1.5B model, where accuracy increases from 15.10% to 35.30%. These findings show that VCSearch has effectively enhanced the model’s ability to model problems.
RQ3. Can VCSearch help existing methods achieve robust mathematical reasoning in realistic scenarios?
In real-world scenarios, mathematical problems rarely fall into strictly well-defined or ill-defined categories. Instead, there is often a need to both solve well-defined problems and identify ill-defined ones. To the best of our knowledge, we are the first to explore this hybrid setting in the context of math word problems (MWP). For our experiments, we employed a balanced sampling strategy to fairly assess the ability to identify ill-posed problems and solve well-defined problems simultaneously. After three repeated experiments, we report the mean ± standard deviation in Table 3.
The results show that VCSearch + CoT and VCSearch + PAL significantly outperform traditional CoT and PAL methods in rejecting unreasonable problems. The rejection rate of ill-defined problems improved by 42.96% and 42.03%, respectively, while the real-world evaluation metrics R-score gained 16.78 and 19.39 points, confirming the application value of the hybrid architecture in complex real-world scenarios.
Table 3: Reaction scores of VCSearch + and comparison methods in a realistic environment with both ill-defined and well-defined problems
Model Methods Reject-Rate R-score
CoT 51.33$\pm$2.29 65.93$\pm$0.73
+Ours 76.13$\pm$1.56 73.98$\pm$0.28
PAL 14.46$\pm$0.41 48.56$\pm$0.22
Qwen2.5 3B +Ours 75.59$\pm$1.39 74.08$\pm$1.17
CoT 39.93$\pm$1.96 53.91$\pm$1.16
+Ours 65.06$\pm$1.48 63.26$\pm$0.84
PAL   7.73$\pm$2.04 32.85$\pm$1.00
Qwen2.5 1.5B +Ours 66.66$\pm$0.24 62.28$\pm$0.65
4.3 More discussion.
Ablations. In this part, we evaluate the impact of two core components of VCSearch on overall performance in Table 4. Removing the iterative search framework(just use one-time refine) results in limited improvement over the baseline SMT solver for few-shot learning. Excluding anchored initialization causes significant search space divergence, with the model becoming overly conservative and rejecting most solutions, severely impairing its ability to solve well-defined data. These findings underscore the necessity of both components.
Table 4: Ablation study on Qwen 7B model.
Search Initialization R-Rate Accuracy
43.59 61.28
89.97 22.81
74.75 77.76
Performance of VCSearch on Models of Different Sizes. Visual analysis of Qwen model results (Figure 4) reveals a strong correlation between model scale and performance: both ill-defined problem identification ability and well-defined problem solving ability decline with smaller models. However, our method mitigates this degradation and even shows advantages across scales. Specifically, VCSearch on Qwen-3B surpasses other methods on Qwen-7B in problem rejection and rivals SMT prompting on models an order of magnitude larger in solving well-defined problems, demonstrating its effectiveness and practical value in resource-limited scenarios.
(a) ill-defined problems
(b) well-defined problems
Figure 4: Performance of VCSearch varying from different model size
5 Related work
Mathematical Reasoning for LLMs. Mathematical reasoning is a crucial aspect in evaluating model reasoning skills, and there are currently two predominant lines for enhancing these skills. One line involves leveraging the existing few-shot prompt tool, such as CoT Wei et al. (2022), PAL Gao et al. (2023). The other is centered around fine-tuning strategy, like Metamath Yu et al. (2023), WizardMath Luo et al. (2023) and Mugglemath Li et al. (2023). Recent work has focused on how to achieve results that match or even exceed those of large models on smaller models Guan et al. (2025) and smaller training datasets Li et al. (2024a) by introducing techniques such as reinforcement learning and MCTS Tolpin and Shimony (2012).
LLM robustness. Previous work can be broadly categorized into two types, perturbations to model inputs and prompting with noisy ground truth. The generation of anticonceptual examples Jia and Liang (2017); Morris et al. (2020); Wang et al. (2021) and irrelevant context Sinha et al. (2019); Clark et al. (2020); Han et al. (2022) are the two main types of input perturbations. In the specific domain of Math Word Problems (MWP),  Zhou et al. (2024) introduces irrelevant punctuation marks as distractors, while  Shi et al. (2023) employs a sentence of unrelated contextual text to serve as a distractor, both aiming to investigate model performance variations. The work most similar to ours is MathTrap Zhao et al. (2024), which focuses on a relatively small set of fewer than 300 ill-defined problems. In contrast, our PMC dataset is far more comprehensive, containing over 5,000 ill-defined problems.
Neuro-Symbolic Methods about LLM reasoning. The primary challenge of these methods lies in ensuring that the LLM correctly translates the reasoning problem from natural language (NL) to the formal language understood by the solver Raza and Milic-Frayling (2025). For instance, Logic-LM Pan et al. (2023) utilizes LLMs to convert natural language into symbolic formulas, SatLM Ye et al. (2024) enables LLMs to generate task specifications that assist in translating natural language into logical predicates. LOT Liu et al. (2024b), similar to CoT, generates progressive logical paths. However, many of these methods struggle to extend successfully to smaller models, due to their limited contextual learning capabilities and lack of formal reasoning knowledge.
6 Conclusion
This paper addresses mathematical reasoning with missing and contradictory conditions by introducing PMC, a large-scale benchmark for evaluating LLM robustness. Our observations reveal a trade-off dilemma between reasoning for well-defined problems and recognizing ill-defined problems. To solve this trade-off, we propose VCSearch, a training-free framework that uses formal language to detect ill-defined problems, enhanced by a variable-constraint pair search strategy to improve formal modeling. Extensive experiments show VCSearch achieves superior robust reasoning across diverse model architectures and sizes.
Limitations
Our work has two main limitations:
Time Consumption. Due to the use of variable-wise refinement and search architecture during the reasoning process, our method incurs higher time overhead compared to the baseline methods.
Limitations of Formal Tools. Our ability to identify ill-defined problems relies on formal tools, such as SMT solver. According to the algorithm design, the system will directly reject tasks that are unsuitable for modeling with logical tools, which may lead to the incorrect rejection of some well-defined problems.
Appendix A Appendix
A.1 Details of PMC
We give more details of our PMC here. We show the number of specific subsets of PMC in Table 5, and show more representative problems to help understand our dataset.
Table 5: The specific number of rewritten datasets
Type AddSub MultiArith SVAMP GSM8k Sum
M-type 369 591 825 1129 2914
C-type 244 359 745 765 2113
Example 1: Example 1 of PMC Statement: Josh decides to try flipping a house. He buys a house for \$80,000 and then puts in \$50,000 in repairs. This increased the value of the house by 150%. How much profit did he make? # Excepted Answer: 70,000 M Version: Josh decides to try flipping a house. He buys a house for \$80,000 and then puts \$50,000 some cost in repairs. This increased the value of the house by 150%. How much profit did he make? C Version: Josh decides to try flipping a house. He buys a house for \$80,000 and then puts in\$50,000 in repairs. This increased the value of the house by 150%, but the market value of the house after repairs is only \$100,000. How much profit did he make? (# market value Contrary to the expected )
Example 2: Example 2 of PMC Statement: Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market? # Excepted Answer: 14 M Version: Janet’s ducks lay 16 eggs per day. She eats three some for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market? C Version: Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg. How much in dollars does she make every day at the farmers’ marketif she give 10 eggs away to her neighbor? (# She only left 9 eggs, can not give away 10 eggs)
A.2 Details of VCSearch
In this part we will introduce the details in our algorithm.
A.2.1 Prompts in VCSearch
We show the prompts we use in VCSearch with examples 3 and 4.
Example 3: prompts used in VCSearch-1 Refine module prompt
I have previously asked you to write Z3 constraints for a problem. However, the current set of constraints for the variable may have omissions or errors. I would like you to review it from the following two aspects and make appropriate modifications if necessary:
1. Based on the problem description, consider whether the current constraints accurately capture the problem. 2. Add constraints based on real-world knowledge, considering whether there are any missing modeling statements, such as the quantity of items should be >= 0, or the relationships between the sides of a triangle. Please note that you only need to add constraints to the CURRENT HEAD VARIABLE; in other words, the new constraints MUST include the head variable! You can first provide your thought process, and then write the new constraints that include the head variable after the identifier <SOS> You can follow the example: Question: Josh decides to try flipping a house. He buys a house for \$80,000 and then puts in \$50,000 in repairs. This increased the value of the house by 150%, but the market value of the house after repairs is only \$100,000. How much profit did he make? Existing Constraints: solver.add(initial-cost == 80000) solver.add(total-investment == initial-cost + repair-cost) solver.add(repair-cost == 50000) solver.add(increased-value-percentage == 0.5) # 150% increase solver.add(expected-value == initial-cost * (100 + increased-value-percentage)) solver.add(market-value-after-repairs == 100000) solver.add(total-investment >= 0) solver.add(profit == market-value-after-repairs - total-investment) solver.add(profit >= 0) solver.add(expected-value >= 0) Now head variable: expected-value Now existing constraints with head variable: solver.add(expected-value == initial-cost * (100 + increased-value-percentage)) solver.add(expected-value >= 0) Answer: 1. For the constraint expected-value == initial-cost * (100 + increased-value-percentage), the equation for expected-value in the problem should be initial-cost * (1 + increased-value-percentage). Therefore, this constraint should be modified to solver.add(expected-value == initial-cost * (1 + increased-value-percentage)). 2. For the constraint solver.add(expected-value >= 0) aligns with real-world requirements. Additionally, since expected-value is an unknown variable, it is appropriate to add real-world constraints, so this should be retained. 3. Furthermore, expected-value and market-value-after-repairs refer to the same entity in the problem, so a constraint should be added: market-value-after-repairs == expected-value. <SOS> So, new Constraints with head variable is solver.add(expected-value == initial-cost * (1 + increased-value-percentage)) solver.add(expected-value >= 0) solver.add(expected-value == market-value-after-repairs) Question:{question} Existing Constraints:{constraint} Now head variable:{head} Now existing constraints with head variable:{constrain-head} Answer:
Example 4: prompts used in VCSearch-2 Verification module prompt
Please judge which set of constraints is better for the given problem, including all constraints of variable "X".
Problem: {question} variable:{head} Constrains set1:{cons1} Constrains set1 ans:{cans1} Constrains set2:{cons2} Constrains set2 ans:{cans1} Please write down your thinking process first, and finally output, "I think Constrains set1 is better", or "I think Constrains set2 is better".
A.2.2 Formal tools
The SMT-LIB(Satisfiability Modulo Theories Library)  Barrett et al. (2010) is a tool for working with satisfiability problems. It provides a standard notation compatible input language for representing logical formulas. And powful SMT solvers, such as Z3 de Moura and Bjørner (2008), extend the classical boolean satisfiability problem (SAT problem) to enable verification of numerical arithmetic problems, among others. The SMT solver will initially determine whether the modeled problem is satisfiable (SAT/UNSAT). If it is satisfiable, the solver will then provide a feasible solution within the feasible domain of the problem. Specifically, we use z3 as a formal tool in the paper.
A.2.3 Double-check solving strategy with SMT solver
We use a double-check strategy when checking with the SMT solver. Specifically, we verify both the satisfiability of the formal expression and the uniqueness of the solution. To be specific, to check the satisfiability of the formal expression, we utilize the Z3 solver. This strategy regards the problem as ill-defined and rejects the answer if the formal expression is unsatisfiable(UNSAT). To assess the uniqueness of the solution, We develop this check through a two-stage process. First, we utilize the Z3 solver to determine one solution and subsequently incorporate this candidate solution as a constraint into the formal expression. If the formal expression remains satisfiable, then it implies that the formal expression encompasses multiple solutions, leading the strategy to reject the answer as it violates the uniqueness of the answer.
To be precise, in the solution phase, our strategy let the SMT solver return four possible different values:
  • Error: Indicates that the modeling cannot be successfully completed. Similar to a compilation error, we do not consider it as a valid state.
  • UNSAT: Indicates that the modeling state cannot be satisfied, there are contradictory conditions, and the answer is rejected.
  • Multi: We believe that the question is ambiguous, resulting in multiple solutions, and the answer is rejected.
  • Ans: Returns a normal real number, representing the answer to the question.
A.3 Details of Experiment
A.3.1 Setup
Compared methods. We selected three representative few-shot prompting methods, along with the zero-shot method that utilizes the intrinsic capabilities of the model, and compared them with our proposed VCSearch method. The methods are introduced as follows: (1)Basic, which is the zero-shot baseline method, directly feeds the problem and instructions to the LLMs without any example problem in the context. (2)CoT,  Wei et al. (2022), requires the model to explicitly output intermediate step-by-step reasoning through natural language before providing the final answer. (3)PAL Gao et al. (2023), converts each step of problem-solving into a programming language format and subsequently utilizes an external programming language interpreter for execution, thereby obtaining the results. (4)SMT, utilizes SMT-LIB to model the problems, then uses an external SMT solver to check for a feasible solution to the problem as well as obtain the ground-truth answer.
Prompts. For the few-shot prompting methods, we prepared four contextual examples (4-shot) for each method, consisting of two well-defined problems and two ill-defined problems. In the system prompt, we explicitly informed the model about the potential presence of ill-defined problems. If the model determines that a problem is unsolvable, it is instructed to output a statement containing the term "unsolvable." This allows us to evaluate whether the model successfully identifies ill-defined problems.
Set up details for Sec4.3. At this part, we employed a balanced sampling strategy to fairly assess the ability to identify ill-posed problems and solve well-defined problems simultaneously. (with a solvable/unsolvable problem ratio of $\alpha=1:1$), selecting 500 samples from the ill-defined problem set (Table 1) and the well-defined problem set (Table 2) to construct a 1000-sample test set. After three repeated experiments, we report the mean ± standard deviation in Table 3.
A.3.2 Prompts used in Preliminary experiments
We show the prompts we use in preliminary experiments to reflect the trade-off dilemma with examples 5.
Example 5: prompts used in Preliminary experiments Pure prompt for ill-defined problem
Now we have some math problems that may be ill-defined. Please judge whether they are indeed ill-defined (no unique real number solution can be determined). If there is indeed no solution, answer true, otherwise answer false. Explain the reason first and then answer.
Pure prompt for well-behaved problem
You’re an experienced elementary school teacher, and I’m now expecting you to solve some math problems.
Mixed prompts
You’re an experienced elementary school teacher, and I’m now expecting you to solve some math problems.  If you find these problems unsolvable, please output “this is unsolvable”. Or please solve this answer, and give the final answer with format "The answer is X"
A.3.3 More experiment results
Table 6: R-Rate on MathTrap
Model Deepcoder Qwen7b Qwen3b Qwen1.5b
Zero 22.95 15.57 15.57 13.72
Ours 65.57 86.06 88.89 74.59
Here, we also tested our method on several other benchmarks that involve refusal to answer. Our method also demonstrated superior performance on MathTrap. However, MathTrap’s mathematical problems involve a significant amount of geometry and algebra, which are not well-suited for formal tool modeling. This is also not suitable for methods such as PAL. So we only compare ours with zero-shot method. In such scenarios, our method adopts a relatively conservative approach, rejecting any problem it cannot confidently solve in order to maintain the safety of the reasoning system.